OpenStudy (anonymous):

The Limit as x ---> infinity of ln(6x + 1) - ln(6x). Please help!!!

OpenStudy (anonymous):

we will use l'hopitals rule here

OpenStudy (anonymous):

but first we need to change it to a proper

OpenStudy (anonymous):

to a proper what? Im not dividing anything here, its subtraction. I thought l'hospitals rule was for rationals.

OpenStudy (anonymous):

are you working on l'hopitals

OpenStudy (anonymous):

no, Im trying to find if a sequence is convergent by taking the limit of the corresponding function.

OpenStudy (anonymous):

But I dont know how to take this limit

OpenStudy (anonymous):

GOT IT

OpenStudy (anonymous):

its ZERO!

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\ln(\frac{6x+1}{6x})=\lim_{x \rightarrow \infty}\ln(1+\frac{1}{6x})\] \[=\ln(\lim_{x \rightarrow \infty}[1+\frac{1}{6x}])=\ln(1+\frac{1}{6}\cdot 0)=\ln(1)=0\]

myininaya (myininaya):

lol ok

OpenStudy (anonymous):

THANK YOU! at least I know im right.

OpenStudy (anonymous):

it is zero from your eyeballs. there is no difference between \[\ln(6x+1)\] and \[\ln(6x)\], at least not for any value of x bigger than ten or so